OpenStudy (pizza114):

The product of 5 and a number x is 1/4 Which statements are true? Select each correct answer. x=120x=120 x+5=14x+5=14 5x=145x=14 x5=14x5=14 x=−194x=−194 x=54

OpenStudy (pizza114):

plz help

OpenStudy (jiteshmeghwal9):

5x=1/4 divide both sides by 5

alones (alones):

Just like @jiteshmeghwal9 saidd You suppose to divide both sides by 5.. \[\frac{ 5x }{ 5 }=\frac{ \frac{ 1 }{ 4 } }{ 5 }\]

OpenStudy (mathmale):

pizza: Please share what you've already done.